OpenStudy (anonymous):

please help :( ...A baseball ht with an initial upward velocity of 100 feet per second will reach height (h) feet after t seconds, as estimated by the formula h-3+100r-16t^2. Assuming nobody touches the baseball, when does it hit the ground? explain your reasoning or show your work round answer to the nearest hundredth

OpenStudy (whpalmer4):

that formula should be \[h = 3+100r-16t^2\]or better yet\[h(t) = 3+100r-16t^2\] What will the height of the ball be when it hits the ground? Can you solve for the value of \(t\) that makes that condition be true?

OpenStudy (whpalmer4):

For super-duper bonus points and a medal for understanding: what is the height of the ball at the instant it is hit?